The wheel is a ring of mass M. Disregard the mas of the spokes. Mass of the wheel is 5 kg. Its radius is 0.5 meters. Angular velocity is 100 rev/min (see the orange arrow for the direction of rotation) Length of axle or shaft is 0.60 meters. a. Calculate the change in the wheel’s angular momentum per second. b. Looking from the top of the setup, is the precession clockwise or counterclockwise? Why?


1
Expert's answer
2021-05-31T15:14:45-0400

Gives

m=5kg


r=0.5meter

Anguler velocity


(w)=100 rev/min

"w=100\times\frac{2\pi}{3\times 60} =\frac{10\pi}{3}Rad/sec"

Saft(l)=0.60 m

Anguler momentum

"J=mvr"


"J=mr^2w"

Pul Value

"J=5\times0.5^2\times {\frac{10\pi }{3}}"

"J=13.08" "Kgm^2\/sec"

Top of see precession is clock wise